# Spacetime Grand Unified Theory

### v1

@@ IV. Consequences of Principle 1
To find the group $G$, we start by noting that the group of rotations $U$ can be divided into two categories: those which affect the vacua $\Omega_{(a)}$ and those that don’t. [^4] Note that $\omega^{\prime\alpha}=U\omega^{\alpha}U^{-1}=\omega^{\alpha}\,,\forall_{\alpha}\Rightarrow U=1_{16\times 16}$, so a similar criteria does not makes sense to apply for $\omega^{\alpha}$. We shall denote the latter as $U_{\omega}$ and the former by $U_{\Omega}$. Focusing on $U_{\omega}$ first, we recall that rotations of $\omega^{\alpha}$ are parametrized by the elements of the spin group Spin(8) Lounesto:2001 . The latter group is a particular example of Spin($2n$), a group of dimension $n(2n-1)$ which can be expressed Mohapatra:2003 as an element of a Clifford algebra $\mathcal{C}\ell_{2n,0}$ of the form
@@
$V^{\prime}_{A}=V_{A}e^{-iH_{\omega}-iY_{\omega}}\,.$ (A176)
@@
which correspond [^5] For each pair $(1,2)$ and $(3,4)$ there is a redundancy in the columns associated to $P_{\perp}$ in each respective quadrant. We could, for example, choose $S_{1}=\textrm{diag}(1,1,1,0,0,0,1,0)$ instead of $S_{1}=\textrm{diag}(0,1,1,0,1,0,1,0)$ and the same change for $S_{2}$, i.e. switch the entries in columns 1 and 5 (respectively entries 12 and 16 for $S_{3}$ and $S_{4}$). One can check that this would be of no consequence for the final results. to a projection into each of the subspaces $\sum_{a}\omega^{\dagger\alpha}\Omega^{\dagger}_{(a)}\Omega_{(a)}\omega^{\alpha}$ for each $\alpha$ plus a fourth component associated to the orthogonal subspace $P_{\perp}$.
@@
where we have performed the rescaling $c_{0}=\theta_{0}/3$ in order for the left-handed electron (corresponding to $V_{12}$) to have the conventional hypercharge $-1$. We thus conclude via Equations (A176), (A195) and (A227) that the group corresponding to the unitary transformation induced by rotations of Witt basis vectors and vacua is precisely the Standard Model group

### v2

@@ IV. Consequences of Principle 1
[…]
@@
$\hskip-8.25128ptV^{\prime}_{A}=V_{A}e^{-iH_{\omega}-iY_{\omega}}\,.$ (A176)
@@
which correspond [^5] For each pair (1,2)(1,2)( 1 , 2 ) and (3,4)(3,4)( 3 , 4 ) there is a redundancy in the columns associated to $P_{\perp}$ in each respective quadrant. We could, for example, choose $S_{1}=\textrm{diag}(1,1,1,0,0,0,1,0)$ instead of $S_{1}=\textrm{diag}(0,1,1,0,1,0,1,0)$ and the same change for $S_{2}$, i.e. switch the entries in columns 1 and 5 (respectively entries 12 and 16 for $S_{3}$ and $S_{4}$). One can check that this would be of no consequence for the final results. to a projection into each of the subspaces $\sum_{a}\omega^{\dagger\alpha}\Omega^{\dagger}_{(a)}\Omega_{(a)}\omega^{\alpha}$ for each $\alpha$ plus a fourth component associated to the orthogonal subspace $P_{\perp}$.
@@
where we have performed the rescaling $c_{0}=\theta_{0}/3$ in order for the left-handed electron (corresponding to $V_{12}$) to have the conventional hypercharge 1-1- 1. We thus conclude via Equations (A176), (A195) and (A227) that the group corresponding to the unitary transformation induced by rotations of Witt basis vectors and vacua is precisely the Standard Model group